Grade 8Algebra

question 21 , rd sharma class 8 , linear equation ex 9.4
plz help me regarding the above question

raghav
5 Years agoGrade 8

1 Answer

Harsh Tiwari
5 Years ago
Let us assume the numerator be x 
The denominator be (2x – 2) 
According formula 
Fraction = numerator/denominator 
= x / (2x – 2) 
So, the numerator and denominator are increased by 3, then fraction is 2/3 
(x + 3)/(2x – 2 + 3) = 2/3 
(x + 3)/(2x + 1) = 2/3 
By doing cross-multiplying 
3(x + 3) = 2(2x + 1) 
3x + 9 = 4x + 2 
3x – 4x = 2 – 9 
-x = -7 
x = 7 
The numerator is x = 7, denominator is (2x – 2) = (2(7) – 2) = 14-2 = 12 
And the fraction is numerator/denominator = 7/12